# CamContextI2V: Context-aware Controllable Video Generation

### v1

@@ Abstract
Recently, image-to-video (I2V) diffusion models have demonstrated impressive scene understanding and generative quality, incorporating image conditions to guide generation. However, these models primarily animate static images without extending beyond their provided context. Introducing additional constraints, such as camera trajectories, can enhance diversity but often degrades visual quality, limiting their applicability for tasks requiring faithful scene representation. We propose CamContextI2V, an I2V model that integrates multiple image conditions with 3D constraints alongside camera control to enrich both global semantics and fine-grained visual details. This enables more coherent and context-aware video generation. Moreover, we motivate the necessity of temporal awareness for an effective context representation. Our comprehensive study on the RealEstate10K dataset demonstrates improvements in visual quality and camera controllability. We make our code and models publicly available at: https://github.com/LDenninger/CamContextI2V.
@@ 1 Introduction
As illustrated in Fig. 1, the initial reference frame alone provides only limited context for the diffusion process. Once the camera pans, the visual quality degrades and arbitrary interpretations of the scene by the diffusion model become evident. To address this, we introduce CamContextI2V, a novel conditioning mechanism that allows users to supply multiple context views, ensuring a comprehensive definition of the scene in which the video is generated.
@@
We propose CamContextI2V, a camera-controllable context-aware diffusion model, which conditions the diffusion process on multiple context frames through a dual-stream encoder retrieving high-level semantic features and low-level visual cues from the context.
@@
Our method achieves a 24.09% improvement in visual quality over the state-of-the-art methods on RealEstate10K dataset.
@@ 4 Method
Image-to-video diffusion models generate videos based on a single reference frame and an optional text condition . Additionally, camera-controlled diffusion models are conditioned on a camera trajectory allowing precise control of the camera view at each timestep. This enforces generations beyond the context conveyed through the reference frame degrading the visual quality. To counteract this, we provide the model with additional context frames and their poses that span a rich context utilized in the diffusion process.
@@ 5 Experiments
| Ours | 53.90 | 45.36 | 2579.96 | 1.53 | 1.09 | 2.29 |
In the following, we thoroughly analyze our method in the setup described in Sec. 5.1. Sec. 5.2 and Sec. 5.3 compares our method to baseline camera-controlled methods and argue the effectiveness of the provided context. Finally, in Sec. 5.4, we investigate our design choices and motivate the importance of 3D and temporal awareness in our method and the complementary effect of our two-stream design. Additionally, we demonstrate the impact of different sampling strategies on our method.
@@ 5.2 Quantitative Comparison
Additionally, we investigate the accuracy of the generated camera trajectory with respect to the RotErr, TransErr and CamMC. We observe a slightly improved rotational error compared to CamI2V’s, indicating an improved camera trajectory of our method. As the evaluation pipeline, GLOMAP, used for estimating the camera trajectory matches keypoint features to simulatenously estimate the camera trajectory and reconstruct a 3D scene using bundle adjustment and we do not train the camera encoder, nor the diffusion model itself, this improved camera trajectory is mainly linked to an improved 3D consistency and visual quality of the generated scene. This demonstrates that the additionally provided context enforces are more faithful representation of the 3D scene.
@@ 5.4 Ablation Studies
Next, we evaluated the effectiveness of our method’s 3D awareness, achieved through the epipolar cross-attention mechanism. Replacing epipolar cross-attention with standard (vanilla) cross-attention, allowing unrestricted feature aggregation from all tokens, still yielded a considerable improvement of FVD points over the baseline. This model variant, still, demonstrates a significant improvement on the baseline by points in the FVD score but fails to match the performance of the 3D-aware model variant. This can be contributed to the model still leveraging the additional context for the generation but failing to reject features from invalid positions, as seen in Fig. 6, especially when context frames provide minimal additional information due to them being sampled from distant regions.
@@
Further, we assess the effect of temporal embeddings integrated into semantic and visual streams. Removing temporal embeddings results in a performance decline, although still outperforming CamI2V considerably. The temporal embeddings, particularly within the visual stream, explicitly guide the temporal attention of the U-Net to properly interpret timestep-specific context. Without this guidance, the epipolar cross-attention timestep-wise embedded context may be interpreted freely, resulting in a impaired performance.
@@
Lastly, Tab. 3 compares different sampling strategies for additional context views. Our default method samples context frames from the interval following the generated video. Further, we investigate two extremes: First, sampling a completely unrelated frame, the furthest frame, as can be seen in Fig. 6. Our results show that this only slightly degrades the visual quality, indicating that our method effectively rejects unrelated features through the induced 3D awareness of the epipolar cross-attention. Second, sampling a frame directly following the video, providing a maximal amount of information to the diffusion process. This only slightly improves our method, showing that it can effectively gather context from loosely placed context views. The qualitative results in Fig. 6 show that our context-aware encoder effectively sorts out unrelated information and provides the diffusion process only with the necessary context.
@@ 6 Conclusion
This paper introduces CamContextI2V, a novel conditioning mechanism that provides the diffusion process with extensive contextual information derived from multiple context views. Unlike conventional image-to-video diffusion models, which typically rely on a single reference image, our proposed method employs a context-aware encoder that encodes additional context through a high-level semantic stream and a 3D-aware visual stream, generating a global semantic representation and a dense, pixel-wise visual embedding from context views. This encoder effectively aggregates relevant features complementing the generation while filtering out unrelated information. As a result, our approach significantly improves visual quality and enhances the accuracy of the generated camera trajectories compared to existing baseline methods.
@@ 7 Additional Qualitative Results
We provide additional qualitative results of our method in Fig. 7.
8 Camera Evaluation
To ensure comparability with the evaluation paradigm proposed by CamI2V, we follow their configuration of the GLOMAP pipeline. The adjusted configuration parameters are listed in Tab. 4. The remaining parameters are set to the default parameters as provided by GLOMAP.
| Stage | Parameter | Value |
| Feature Extraction | ImageReader.single_camera | 1 |
| ImageReader.camera_model | SIMPLE_PINHOLE | |
| ImageReader.camera_params | ||
| SiftExtraction.estimate_affine_shape | 1 | |
| SiftExtraction.domain_size_pooling | 1 | |
| Sequential Matcher | SiftMatching.guided_matching | 1 |
| SiftMatching.max_num_matches | 65536 | |
| Mapper | RelPoseEstimation.max_epipolar_error | 4 |
| BundleAdjustment.optimize_intrinsics | 0 |

### v2

# CamC2V: Context-aware Controllable Video Generation

@@ Abstract
Recently, image-to-video (I2V) diffusion models have demonstrated impressive scene understanding and generative quality, incorporating image conditions to guide generation. However, these models primarily animate static images without extending beyond their provided context. Introducing additional constraints, such as camera trajectories, can enhance diversity but often degrade visual quality, limiting their applicability for tasks requiring faithful scene representation. We propose CamC2V, a context-to-video (C2V) model that integrates multiple image conditions as context with 3D constraints alongside camera control to enrich both global semantics and fine-grained visual details. This enables more coherent and context-aware video generation. Moreover, we motivate the necessity of temporal awareness for an effective context representation. Our comprehensive study on the RealEstate10K dataset demonstrates improvements in visual quality and camera controllability. We will publish our code upon acceptance.
@@ 1 Introduction
As illustrated in Fig. 1, the initial reference frame alone provides only limited context for the diffusion process. Once the camera pans, the visual quality degrades and arbitrary interpretations of the scene by the diffusion model become evident. To address this, we introduce CamC2V, a novel conditioning mechanism that allows users to supply multiple context views, ensuring a comprehensive definition of the scene in which the video is generated.
@@
We propose CamC2V, a camera-controllable context-aware diffusion model, which conditions the diffusion process on multiple context frames through a dual-stream encoder retrieving high-level semantic features and low-level visual cues from the context.
@@
Our method achieves a 24.09% improvement in visual quality over the state-of-the-art methods on the RealEstate10K dataset.
@@ 4 Method
Image-to-video diffusion models generate videos based on a single reference frame and an optional text condition . Additionally, camera-controlled diffusion models are conditioned on a camera trajectory allowing precise control of the camera view at each timestep. The reference frame does not always provide the necessary context corresponding to the camera trajectory. This can lead to insufficient visual quality of the generated frames. In contrast, we propose a new scheme coined context-to-video which enhances the generation process with a rich context conveyed through additional context frames and their poses
@@
Log-weighted Loss.
Our Context-aware Encoder injects crucial information, especially to later frames, that have limited context in the baseline methods leading to degrading visual quality. To force the training process to focus on such frames that rely most heavily on our context‐aware conditioning, we apply a logarithmic re‐weighting to the standard reconstruction loss along the time axis. Specifically, for each frame in the sequence, we define:
This not only improves generative quality but also stabilizes the training, mitigating divergence in later stages of training.
@@ 5 Experiments
[…]
In the following, we thoroughly analyze our method in the setup described in Sec. 5.1. Sec. 5.2 and Sec. 5.3 compares our method to baseline camera-controlled methods and argue the effectiveness of the provided context. Next, in Sec. 5.4, we investigate our design choices and motivate the importance of 3D and temporal awareness in our method and the complementary effect of our two-stream design. Finally, in Sec. 5.5 we compare our method to a state-of-the-art Novel View Synthesis approach, employing similar conditioning signals.
@@ 5.2 Quantitative Comparison
Additionally, we investigate the accuracy of the generated camera trajectory with respect to the RotErr, TransErr and CamMC. We observe a slightly improved rotational error compared to CamI2V’s, indicating an improved camera trajectory of our method. As the evaluation pipeline, GLOMAP, used for estimating the camera trajectory matches keypoint features to simultaneously estimate the camera trajectory and reconstruct a 3D scene using bundle adjustment and we do not train the camera encoder, nor the diffusion model itself, this improved camera trajectory is mainly linked to an improved 3D consistency and visual quality of the generated scene. This demonstrates that the additionally provided context enforces are more faithful representation of the 3D scene.
@@ 5.4 Ablation Studies
Next, we evaluated the effectiveness of our method’s 3D awareness, achieved through the epipolar cross-attention mechanism. Replacing epipolar cross-attention with standard (vanilla) cross-attention, allowing unrestricted feature aggregation from all tokens, still yielded a considerable improvement of FVD points over the baseline. This model variant, still, demonstrates a significant improvement on the baseline by points in the FVD score but fails to match the performance of the 3D-aware model variant. This can be attributed to the model still leveraging the additional context for the generation but failing to reject features from invalid positions, as seen in Fig. 6, especially when context frames provide minimal additional information due to them being sampled from distant regions.
@@
Further, we assess the effect of temporal embeddings integrated into semantic and visual streams. Removing temporal embeddings results in a performance decline, although still outperforming CamI2V considerably. The temporal embeddings, particularly within the visual stream, explicitly guide the temporal attention of the U-Net to properly interpret timestep-specific context. Without this guidance, the epipolar cross-attention timestep-wise embedded context may be interpreted freely, resulting in impaired performance.
@@
Lastly, Tab. 3 compares different sampling strategies for additional context views. Our default method samples context frames from the interval following the generated video. Furthermore, we investigate two extremes: first, sampling a completely unrelated frame, the furthest frame, as shown in Fig. 6. Our results show that this only slightly degrades the visual quality, indicating that our method effectively rejects unrelated features through the induced 3D awareness of the epipolar cross-attention. Second, sampling a frame directly following the video, providing a maximal amount of information to the diffusion process. This only slightly improves our method, showing that it can effectively gather context from loosely placed context views. The qualitative results in Fig. 6 show that our context-aware encoder effectively sorts out unrelated information and provides the diffusion process only with the necessary context.
5.5 Comparison to Novel View Synthesis Approach
We finally compare our method against FrugalNeRF [9], a state-of-the-art novel view synthesis approach for sparse views. Such NeRF-based methods typically rely on an initial sparse 3D reconstruction step, limiting their applicability to diverse scenes. In a two-view setup, the COLMAP pipeline [16], typically employed in this step, only achieves a reliable registration in over of the cases if the frames distance is between and , heavily limiting such approaches if the context frames are too close or too distant.
We train FrugalNeRF on test scenes from the RealEstate10K dataset using the first and th frame as training frames and the intermediate ones as test frames, similar to the End+1 evaluation setup of the context sampling study in Sec. 5.4. Accordingly, we define the reference and context frame for our approach.
Tab. 4 shows that our method slightly outperforms FrugalNeRF while being more broadly applicable, as it does not require a prior reconstruction step. Moreover, because NeRF-based methods typically employ a test-time optimization training scheme, the time needed to learn and render the 3D representation () is substantially larger than the pure rendering time for our approach (), which can be directly applied to novel scenes after training. See the supplementary material for additional details.
| Method | SSIM | LPIPS | PSNR | Time (s) |
| FrugalNeRF | 0.737 | 0.156 | 23.76 | 1265.87 |
| Ours | 0.741 | 0.128 | 24.32 | 8.01 |
@@ 6 Conclusion
This paper introduces CamC2V, a novel conditioning mechanism that provides the diffusion process with extensive contextual information derived from multiple context views. Unlike conventional image-to-video diffusion models, which typically rely on a single reference image, our proposed method employs a Context-aware Encoder that encodes additional context through a high-level semantic stream and a 3D-aware visual stream, generating a global semantic representation and a dense, pixel-wise visual embedding from context views. This encoder effectively aggregates relevant features complementing the generation while filtering out unrelated information. As a result, our approach significantly improves visual quality and enhances the accuracy of the generated camera trajectories compared to existing baseline methods.
@@
7 Computational Demand
Table 5 reports peak GPU memory and end-to-end inference latency for the baseline models and for our method. As shown, the Context-aware Encoder introduces only a small computational overhead relative to the base model. The visual stream exhibits a larger (yet still minor) latency than the semantic stream because it employs additional timestep- and pixel-wise query tokens to produce a dense visual representation.
| Method | # Params | GPU Memory (GiB) | Latency (s) | |
| Trainable | Total | |||
| DynamiCrafter | - | 2.6 B | 10.43 | 5.143 |
| MotionCtrl | - | 2.6 B | 10.49 | 4.849 |
| CameraCtrl | - | 2.8 B | 11.28 | 4.898 |
| CamI2V | - | 2.9 B | 11.21 | 7.976 |
| CamC2V | 97.4 M | 2.9 B | 11.68 | 8.208 |
| – Semantic Stream | 50.9 M | 50.9 M | 0.19 | 0.004 |
| – Visual Stream | 46.5 M | 46.5 M | 0.17 | 0.043 |
8 Camera Evaluation
We assess the quality of the generated camera trajectories by estimating camera poses with the GLOMAP pipeline. The non-default configuration used in our evaluation is listed in Table 7.
| Parameter | Value |
| ImageReader.single_camera | 1 |
| ImageReader.camera_model | SIMPLE_PINHOLE |
| ImageReader.camera_params | |
| SiftExtraction.estimate_affine_shape | 1 |
| SiftExtraction.domain_size_pooling | 1 |
| SiftMatching.guided_matching | 1 |
| SiftMatching.max_num_matches | 65536 |
| RelPoseEstimation.max_epipolar_error | 4 |
| BundleAdjustment.optimize_intrinsics | 0 |
@@ 9 Additional Qualitative Results
To further assess the generative quality we provide further qualitative results of our method in Fig. 9 compared to the baseline method CamI2V.
10 COLMAP Evaluation
NeRF or 3D Gaussian Splatting-based approaches often rely on an initial sparse 3D reconstruction step to get an initial pointcloud of the scene. Typically, COLMAP [16] is used as a sparse 3D reconstruction pipeline. The pipeline can be split in three stages, namely feature extraction, feature matching and point triangulation. After the first stage, we insert the ground-truth camera poses into the pipeline to improve the triangulation performance and kepp it fair to our model, which is also provided with ground-truth camera poses.
| Parameter | Value |
| ImageReader.single_camera | 1 |
| ImageReader.camera_model | SIMPLE_PINHOLE |
| ImageReader.camera_params | |
| Mapper.tri_ignore_two_view_tracks | 0 |
| Mapper.num_threads | 16 |
| Mapper.init_min_tri_angle | 1 |
| Mapper.multiple_models | 0 |
| Mapper.extract_colors | 0 |
11 Novel View Synthesis Qualitative Results
We provide additional qualitative results comparing our model against FrugalNeRF in Fig. 9. It is visible that FrugalNeRF not only suffers from an insufficient initial sparse reconstruction step, indicated by the black borders, but also introduces artifacts, especially for fine-grained details and edges, where the 3D representation is not sufficiently learned. Overall, our approach mitigates such artifacts and produces clearer more realistic looking images.
In Fig. 8, we show the success rate of this pipeline plotted against the distance between the two used frames. All failure cases come from the triangulation stage, where the model fails to confidently match and triangulate points between images. Especially, for larger frame distances the success rate significantly drops, eventhough the images show sufficient overlap in our visual inspection. This ultimately only leaves a small window from to where the pipeline is able to reliably produce a sparse 3D reconstruction in over of the cases. Overall, in our used evaluation setup using the first and th frame the pipeline has a success rate of .